OpenStudy (xgamer7868):

will medal and fan help asap i got 10 mins to do this! You and your friend are splitting a restaurant bill and leaving a tip. You pay half of the bill plus $5. Your friend pays half of the bill plus 25% of half the bill. You and your friend pay the same total amount. How much is half the bill? (a) Write an equation. Let x be the cost of half the bill. (b) Solve the equation. Show your work. (c) Check your solution. Show your work. (d) State the solution in complete sentences.

OpenStudy (xgamer7868):

please help

OpenStudy (halloweeninheaven):

a = Your Payment b = Your Friend's Payment x = Bill What you're trying to find = 12x12x+5=a 12x+25% of 12x=b 25%=14 Since you and your friend pay the same total, a=b If a=b,12x+5=12x+14(12x) First, you want to simplify. 12x+5=12x+18x Multiply by 8 on both sides to get rid of those fractions 4x+40=4x+x Subtract 4x on both sides to find x 40=x you're looking for HALF the bill, so it'd be 12 of x Answer = 20

OpenStudy (xgamer7868):

that all i need?

OpenStudy (halloweeninheaven):

yes i had this question used this and passed

OpenStudy (xgamer7868):

k

OpenStudy (xgamer7868):

can u help with one more?

OpenStudy (halloweeninheaven):

i can try what is the question?

OpenStudy (xgamer7868):

ill paste it hold on

OpenStudy (halloweeninheaven):

ok

OpenStudy (xgamer7868):

cant past

OpenStudy (xgamer7868):

well i gtg

OpenStudy (xgamer7868):

thx

OpenStudy (halloweeninheaven):

yeah no problem